OpenStudy (anonymous):

Karen and Carolyn take turns mowing the lawn. Karen mowed 3/8 of the lawn of the first day and on the 2nd day Carolyn mowed 1/3 of what was left to be mowed. What fraction of the lawn remains to be mowed after the 2nd day?

OpenStudy (anonymous):

1 - 3/8 = 5/8 (5/8) x 1/3 = 5/24 5/ 8 - 5/24 = 10/24 = 5/12